# SDEval: Safety Dynamic Evaluation for Multimodal Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) (Brown et al. 2020; Achiam et al. 2023; Reid et al. 2024; DeepSeek-AI et al. 2024) have achieved significant advancements. Recent developments have extended this success into the multi-modal realm, allowing LLMs to execute various high-level vision tasks, including visual content understanding and generation (Xie et al. 2024; Li et al. 2024b; Gao et al. 2025; OpenAI 2024, 2025; Zhao et al. 2025). Despite the success in MLLM capabilities, there is a huge risk that MLLMs may generate outputs that diverge from their creators’ intended goals, potentially resulting in untruthful or harmful content (Hendrycks, Mazeika, and Woodside 2023; Yao et al. 2024). This highlights the crucial need for ensuring MLLM safety before deployment. Comprehensive assessment of their potential risks and corresponding mitigation strategies are needed.
Recently, several studies have initiated preliminary explorations into evaluating the safety of MLLMs. MLLMGuard (Gu et al. 2024) provides the safety analysis in both English and Chinese, using data from social media. Hu et al. (2024) identified information leakage issues in the existing datasets and proposed VLSBench, improving evaluation accuracy by better aligning image and text modalities. Besides, existing efforts also establish relatively comprehensive safety evaluation systems (Ying et al. 2024; Zhang et al. 2024; Cai et al. 2024; Liu et al. 2023b). However, after reviewing existing benchmarks, we identify the following main challenges in achieving reliable safety evaluation: 1) Data leakage. Most safety benchmarks build their dataset by integrating open-source datasets(Zhang et al. 2024; Gu et al. 2024; Zhang et al. 2025a; Xia et al. 2025; Liu et al. 2025), which are likely to be included in the MLLM training sets. Affected by this, the results of MLLMs on these benchmarks may lead to concerns, causing a misunderstanding in the entire community. 2) Static dataset with fixed complexity. Existing MLLM safety benchmarks are manually constructed and lack updating. Their fixed complexity can’t match the fast progress of MLLM. To gauge MLLM performance limits precisely, there’s an urgent need for a dynamic, automated evaluation framework with adjustable complexity. 3) Attack methods continue to evolve. As new attack methods emerge, MLLM safety benchmarks should be updated accordingly to further test model safety performance. Although previous studies have proposed simple dynamic evaluation methods (Yang et al. 2024b; Zhu et al. 2023), they are typically applied to model capability evaluation, neglecting safety and the balance of capability-safety, which hinders effective dynamic safety evaluation.
@@
We leverage SDEval for representative safety evaluation benchmarks such as MLLMGuard (Gu et al. 2024) and VLSBench (Hu et al. 2024), and capability evaluation benchmarks, MMVet (Yu et al. 2023) and MMBench (Liu et al. 2023d). Experiments on various MLLMs, e.g., GPT-4o (OpenAI 2024), Claude-4-Sonnet (Anthropic. 2025), and DeepSeek-VL family (Lu et al. 2024), demonstrate that SDEval impacts the safety of different MLLMs to varying degrees, with InternVL-3-78B experiencing a safety reduction of nearly 10%. These results indicate that our dynamic strategy significantly alleviates the data leakage problem, changes data distribution, and increases dataset complexity.
@@ 2.1 Data Contamination
MLLMs are often pre-trained on massive, diverse datasets—often scraped from the web, which increases the risk of evaluation data overlapping (Dodge et al. 2021; Zhou et al. 2023). In the post-training phase, models are further fine-tuned on large human-annotated or synthetic datasets that may resemble evaluation tasks, further compounding contamination risks. Although retrieval-based detection methods (Golchin and Surdeanu 2023a, b; Yang et al. 2023) exist, the sheer scale and complexity of training corpora make it difficult to entirely exclude evaluation data. Additionally, many MLLMs keep their training data proprietary, complicating the accurate assessment of their true performance and highlighting the need for fair and reliable benchmarks. To address this issue, dynamic benchmarking has been proposed (Zhu et al. 2023; Fan et al. 2023; Lei et al. 2023). In this paper, we make a step forward in the safety dynamic benchmarking for MLLMs.
@@ 2.2 MLLM Safety Evaluation
Despite the great success of MLLMs in multimodal understanding and reasoning, their potential safety issues, such as truthfulness, value misalignment, and misuse, still pose significant challenges. Efforts have been made to evaluate the safety of MLLMs. Liu et al. (2023b) proposed MMsafetybench, a VQA dataset covering 13 harmful scenarios to assess MLLMs’ safety. Ch3ef (Shi et al. 2024) adopts ”Helpful, Honest, and Harmless” as safety evaluation criteria. Other benchmarks Hu et al. (2024); Gu et al. (2024) also investigate safety from different degrees. However, these benchmarks are manually constructed and lack updating. Their fixed complexity and diversity can’t match the fast progress of MLLMs. To address this issue, we propose SDEval to make dynamic evaluation, which injects randomness into existing data for MLLM safety benchmarking.
@@ 2.3 Dynamic Evaluation
Recently, the evaluations for MLLMs have gained much attention from both academia and industry (Yao 2025). However, existing evaluation benchmarks are static and have data leakage issues, so they can not match the fast progress of MLLMs. To solve this, some researchers have pioneered the exploration of dynamic evaluation. Zhu et al. propose DyVal V1 and V2 for the evaluation of LLMs (Zhu et al. 2023, 2024). Recently, Yang et al. (2024b) and Zhou et al. (2025) transfer this insight into the multi-modal domain and propose a vision-language bootstrapping strategy. However, these methods are not suitable for safety evaluation due to that their scenarios are limited to specific multiple-choice questions, while safety evaluation is usually open-ended and does not have fixed answers. Different from existing strategies, we construct a safety-centric dynamic framework, SDEval. Starting from any original benchmark, we can endlessly generate the variants with flexible complexity and lower contamination rates.
@@ 3.2 Text Dynamics
Language understanding is crucial for MLLMs, recent jailbreak research of MLLMs (Chao et al. 2023; Deng et al. 2023; Liu et al. 2023c) revealed that current MLLMs are sensitive to input texts. Similar to DME (Yang et al. 2024b), we construct text dynamics from the human-centric perspective–humans often adopt strategies such as replacing sensitive words, reorganizing sentences, or combining multiple languages (e.g., English and Chinese in one sentence) to form new sentences while maintaining semantics to circumvent safety review. Specifically, we utilize six dynamic strategies to modify the text prompt of sample :
@@
Given that some safety review mechanisms usually identify keywords, it is effective and reasonable to replace the words in the original sentence to perform text dynamic processing. Inspired by (Zhu et al. 2024), we prompt LLMs to replace no more than five words of each text prompt, , using synonyms or contextually similar words. For example, the word religious may be replaced by its synonyms like faith-based or faithful.
@@
Following DME (Yang et al. 2024b), we utilize GPT-4o (OpenAI 2024) to add extra relevant/irrelevant descriptions into the original text, which may distract the model’s attention, thereby reducing its control over safety. Specifically, for adding relevant descriptions, we employ GPT-4o (OpenAI 2024) to analyze the image and generate a caption about the image, then add the caption in front of the original text; for irrelevant descriptions, we prompt GPT-4o (OpenAI 2024) to add descriptions that are not related to the image , and we append the irrelevant descriptions after the original text prompt.
@@
Given the fact that humans often deliberately make spelling mistakes or repeat certain letters in specific words to evade safety review. These operations will not change the meaning of the sentence and will not affect the reader’s normal reading. Similar to (Vega et al. 2024), we utilize GPT-4o to make typos for each word in the given sentence by randomly selecting from the strategies of repeating, spelling mistakes, and special wrong characters.
@@
Considering the potential safety risks caused by the inaccurate recognition of multiple languages, we use a multilingual hybrid strategy to dynamically update text . Specifically, we prompt GPT-4o (OpenAI 2024) to reconstruct the original sentence from a single language into a combination of multiple languages, including Chinese, English, Russian, French, Japanese, and Korean.
@@ 3.3 Image Dynamics
Harmfulness in image content is also critical for MLLM safety evaluation. Pioneering research (Yang et al. 2024b) proves that significant overlap exists between evaluation benchmark and MLLM training data, leading to data leakage problems. To address this, we adopt the strategy of dynamically updating and evaluating. The image strategies utilized can be divided into two parts:
@@
1) Spatial Transformation: Inspired by (Zhou et al. 2025), we expand the original image boundaries by random padding, whose width is uniformly sampled from the range of the original image size. After that, random flipping is also adopted. This operation helps in testing MLLMs’ ability to recognize and locate harmful objects under different spatial transforms.
@@
1) For generation, we prompt GPT-4o (OpenAI 2024) to generate a comprehensive caption, which highlights the layout, object details, and scene descriptions. Then, we guide Stable-Diffusion-3.5-Large (Esser et al. 2024) with the caption to generate a new version of the original image. Then, we leverage GPT-4o (OpenAI 2024) to verify whether the generated image maintains essential concepts of the original one. If not, we repeat this process until success.
2) For manipulation, we edit the original image via inserting objects, inserting texts, and style transfer. Specifically, we prompt GPT-4o (OpenAI 2024) to analyze the image and answer whether and how to conduct the following editing: (a) Inserting objects: Return the name of an object that can be inserted into the image without affecting the main content and safety of the image and the location that the object should be inserted; (b) Inserting texts: Return appropriate texts that can be inserted into the image without affecting semantics of the image. (c) Style transfer: Choose a suitable style from {Watercolor style, Sketch style, Comic style} based on the image. Then we conduct manipulation operations by utilizing the ICEdit (Zhang et al. 2025c) model. Examples are presented in Figure 4.
@@ 3.4 Text-Image Dynamics
Section 3.2 introduces a set of text dynamic strategies. In this text-to-image strategy, we aim to inject text perturbations into images via cross-modal generation, thereby impacting MLLM safety. This strategy amplifies the influence of text dynamics through cross-modal interactions. Specifically, we first sample a from text dynamics, then we feed the sampled and original image into GPT-4o to generate an image caption and extract the safety-related keywords, which will then be used as a prompt to generate new images via stable diffusion (Esser et al. 2024). We then utilize the sampled text and the generated image as the new text-image pair.
@@
We mainly investigate two jailbreaking tricks. a) Figstep (Gong et al. 2023) proves that MLLMs can recognize and answer the typographic questions in images. The safety guardrails of MLLMs are ineffective against the typographic visual prompts. Even if the LLM part of MLLMs has been safety aligned in advance, the visual inputs could introduce new risks since the visual embedding space is not safety aligned to the LLM’s embedding space. Based on this, we directly replace the text prompts with their typographic version. Thus, in this trick, only the typographic prompt and the original image are fed into MLLMs. b) HADES (Li et al. 2024a) empirically validates that MLLMs can be significantly affected by the unsafe words contained in images rather than texts, so we utilize GPT-4o to extract key information, which is strongly related to the safety content, and then inject it into the original image. This trick moves the unsafe content from the texts to the images without changing the original semantics.
@@ 4.1 Experimental Setup
In order to evaluate the effectiveness of SDEval, we select two comprehensive multimodal large language models safety benchmark MLLMGuard (Gu et al. 2024) and VLSBench (Hu et al. 2024) for dynamic evaluation.
MLLMguard (Gu et al. 2024) develops a collection of adversarial examples to test the ability of MLLMs to identify and counteract attacks orchestrated by red teams. We follow (Gu et al. 2024) to evaluate the proposed two metrics, Attack Success Degree (ASD), which quantifies the extent to which the responses are harmless, and Perfect Answer Rate (PAR), which represents the proportion of safe and responsible responses among all responses.
VLSBench (Hu et al. 2024) aims to address the existing issues in current multimodal safety benchmarks, called Visual Safety Information Leakage (VSIL). Following VLSBench, we calculate the Safety Rate (SR) by considering the total number of safe refusals and safe warnings.
@@
We evaluate four close source MLLMs: GPT-4o (OpenAI 2024), o3 (OpenAI 2025), Claude-4-Sonnet (Anthropic. 2025), and Gemini2.5-Pro (Gemini Team 2025), and extensive open-sourced models: Qwen-VL family (Bai et al. 2025): Qwen2.5-VL-7B, Qwen2.5-VL-72B, Qwen2-VL-7B, Qwen2-VL-72B; Yi-VL family (Young et al. 2024): Yi-VI-6B, Yi-VI-34B; InternVL family (Zhu et al. 2025): InternVL-3-9B, InternVL-3-14B, InternVL-Chat-V1-5; LLaVA family (Liu et al. 2023a): LLaVA-V1.5-7B, LLaVA-V1.5-13B. To ensure a standardized comparison, we set the generation temperature to 0 for all models. All experiments are conducted on 4 NVIDIA A800 GPUs.

### v2

@@ 1 Introduction
Large language models (LLMs) (Achiam2023GPT4TR; Reid2024Gemini1U) have achieved significant advancements. Recent developments have extended this success into the multi-modal realm, allowing LLMs to execute various high-level vision tasks, including visual content understanding and generation (Xie2024ShowoOS; Li2024HunyuanDiTAP; Gao2025Seedream3T; gpt4o; openai2025o3; Zhao2025CanPI; wang2025affordance; yuan2024cultural; zhang2025xiongkuo). Despite the success in MLLM capabilities, there is a huge risk that MLLMs may generate outputs that diverge from their creators’ intended goals, potentially resulting in untruthful or harmful content (Hendrycks2023AnOO; YAO2024100211; zhang2025lmmsurvey; zhu2025safetyflow). This highlights the crucial need for ensuring MLLM safety before deployment. Comprehensive assessment of their potential risks and corresponding mitigation strategies are needed.
Recently, several studies have initiated preliminary explorations into evaluating the safety of MLLMs. MLLMGuard (gu2024mllmguard) provides the safety analysis in both English and Chinese, using data from social media. hu2024vlsbench identified information leakage issues in the existing datasets and proposed VLSBench, improving evaluation accuracy by better aligning image and text modalities. Besides, existing efforts also establish relatively comprehensive safety evaluation systems (ying2024safebench; zhang2024multitrust; cai2024benchlmm; Liu2023MMSafetyBenchAB). However, after reviewing existing benchmarks, we identify the following main challenges in achieving reliable safety evaluation: 1) Data leakage. Most safety benchmarks build their dataset by integrating open-source datasets(zhang2024multitrust; gu2024mllmguard; zhang2025spa; xia2025msr; liu2025guardreasoner), which are likely to be included in the MLLM training sets. Affected by this, the results of MLLMs on these benchmarks may lead to concerns, causing a misunderstanding in the entire community. 2) Static dataset with fixed complexity. Existing MLLM safety benchmarks are manually constructed and lack updating. Their fixed complexity can’t match the fast progress of MLLM. To gauge MLLM performance limits precisely, there’s an urgent need for a dynamic, automated evaluation framework with adjustable complexity. 3) Attack methods continue to evolve. As new attack methods emerge, MLLM safety benchmarks should be updated accordingly to further test model safety performance. Although previous studies have proposed simple dynamic evaluation methods (Yang2024DynamicME; Zhu2023DyValDE), they are typically applied to model capability evaluation, neglecting safety and the balance of capability-safety, which hinders effective dynamic safety evaluation.
@@
We leverage SDEval for representative safety evaluation benchmarks such as MLLMGuard (gu2024mllmguard) and VLSBench (hu2024vlsbench), and capability evaluation benchmarks, MMVet (yu2023mm) and MMBench (Liu2023MMBenchIY). Experiments on various MLLMs, e.g., GPT-4o (gpt4o), Claude-4-Sonnet (claude_4_opus), and DeepSeek-VL family (lu2024deepseekvl), demonstrate that SDEval impacts the safety of different MLLMs to varying degrees, with InternVL-3-78B experiencing a safety reduction of nearly 10%. These results indicate that our dynamic strategy significantly alleviates the data leakage problem, changes data distribution, and increases dataset complexity.
@@ 2.1 Data Contamination
MLLMs are often pre-trained on massive, diverse datasets—often scraped from the web, which increases the risk of evaluation data overlapping (Dodge2021DocumentingLW; Zhou2023DontMY). In the post-training phase, models are further fine-tuned on large human-annotated or synthetic datasets that may resemble evaluation tasks, further compounding contamination risks. Although retrieval-based detection methods (Golchin2023TimeTI; Yang2023RethinkingBA) exist, the sheer scale and complexity of training corpora make it difficult to entirely exclude evaluation data. Additionally, many MLLMs keep their training data proprietary, complicating the accurate assessment of their true performance and highlighting the need for fair and reliable benchmarks. To address this issue, dynamic benchmarking has been proposed (Zhu2023DyValDE; fan2023nphardeval; lei2023s3eval). In this paper, we make a step forward in the safety dynamic benchmarking for MLLMs.
@@ 2.2 MLLM Safety Evaluation
Despite the great success of MLLMs in multimodal understanding and reasoning, their potential safety issues, such as truthfulness, value misalignment, and misuse, still pose significant challenges. Efforts have been made to evaluate the safety of MLLMs. Liu2023MMSafetyBenchAB proposed MMsafetybench, a VQA dataset covering 13 harmful scenarios to assess MLLMs’ safety. Ch3ef (Shi2024AssessmentOM) adopts ”Helpful, Honest, and Harmless” as safety evaluation criteria. Other benchmarks hu2024vlsbench; gu2024mllmguard also investigate safety from different degrees. However, these benchmarks are manually constructed and lack updating. Their fixed complexity and diversity can’t match the fast progress of MLLMs. To address this issue, we propose SDEval to make dynamic evaluation, which injects randomness into existing data for MLLM safety benchmarking.
@@ 2.3 Dynamic Evaluation
Recently, the evaluations for MLLMs have gained much attention from both academia and industry (NextAI; tian2024coding; tian2025smc++; tian2025rofi; tian2025towards; tian2025semantic). However, existing evaluation benchmarks are static and have data leakage issues, so they can not match the fast progress of MLLMs. To solve this, some researchers have pioneered the exploration of dynamic evaluation. Zhu et al. propose DyVal V1 and V2 for the evaluation of LLMs (Zhu2023DyValDE; Zhu2024DyVal2D). Recently, Yang2024DynamicME and Zhou2025MDK12BenchAM transfer this insight into the multi-modal domain and propose a vision-language bootstrapping strategy. However, these methods are not suitable for safety evaluation due to that their scenarios are limited to specific multiple-choice questions, while safety evaluation is usually open-ended and does not have fixed answers. Different from existing strategies, we construct a safety-centric dynamic framework, SDEval. Starting from any original benchmark, we can endlessly generate the variants with flexible complexity and lower contamination rates.
@@ 3.2 Text Dynamics
Language understanding is crucial for MLLMs, recent jailbreak research of MLLMs (Chao2023JailbreakingBB; Deng2023MultilingualJC; Liu2023JailbreakingCV) revealed that current MLLMs are sensitive to input texts. Similar to DME (Yang2024DynamicME), we construct text dynamics from the human-centric perspective–humans often adopt strategies such as replacing sensitive words, reorganizing sentences, or combining multiple languages (e.g., English and Chinese in one sentence) to form new sentences while maintaining semantics to circumvent safety review. Specifically, we utilize six dynamic strategies to modify the text prompt of sample :
@@
Given that some safety review mechanisms usually identify keywords, it is effective and reasonable to replace the words in the original sentence to perform text dynamic processing. Inspired by (Zhu2024DyVal2D), we prompt LLMs to replace no more than five words of each text prompt, , using synonyms or contextually similar words. For example, the word religious may be replaced by its synonyms like faith-based or faithful.
@@
Following DME (Yang2024DynamicME), we utilize GPT-4o (gpt4o) to add extra relevant/irrelevant descriptions into the original text, which may distract the model’s attention, thereby reducing its control over safety. Specifically, for adding relevant descriptions, we employ GPT-4o (gpt4o) to analyze the image and generate a caption about the image, then add the caption in front of the original text; for irrelevant descriptions, we prompt GPT-4o (gpt4o) to add descriptions that are not related to the image , and we append the irrelevant descriptions after the original text prompt.
@@
Given the fact that humans often deliberately make spelling mistakes or repeat certain letters in specific words to evade safety review. These operations will not change the meaning of the sentence and will not affect the reader’s normal reading. Similar to (vega2024stochastic), we utilize GPT-4o to make typos for each word in the given sentence by randomly selecting from the strategies of repeating, spelling mistakes, and special wrong characters.
@@
Considering the potential safety risks caused by the inaccurate recognition of multiple languages, we use a multilingual hybrid strategy to dynamically update text . Specifically, we prompt GPT-4o (gpt4o) to reconstruct the original sentence from a single language into a combination of multiple languages, including Chinese, English, Russian, French, Japanese, and Korean.
@@ 3.3 Image Dynamics
Harmfulness in image content is also critical for MLLM safety evaluation. Pioneering research (Yang2024DynamicME) proves that significant overlap exists between evaluation benchmark and MLLM training data, leading to data leakage problems. To address this, we adopt the strategy of dynamically updating and evaluating. The image strategies utilized can be divided into two parts:
@@
1) Spatial Transformation: Inspired by (Zhou2025MDK12BenchAM), we expand the original image boundaries by random padding, whose width is uniformly sampled from the range of the original image size. After that, random flipping is also adopted. This operation helps in testing MLLMs’ ability to recognize and locate harmful objects under different spatial transforms.
@@
1) For generation, we prompt GPT-4o (gpt4o) to generate a comprehensive caption, which highlights the layout, object details, and scene descriptions. Then, we guide Stable-Diffusion-3.5-Large (Esser2024ScalingRF) with the caption to generate a new version of the original image. Then, we leverage GPT-4o (gpt4o) to verify whether the generated image maintains essential concepts of the original one. If not, we repeat this process until success.
2) For manipulation, we edit the original image via inserting objects, inserting texts, and style transfer. Specifically, we prompt GPT-4o (gpt4o) to analyze the image and answer whether and how to conduct the following editing: (a) Inserting objects: Return the name of an object that can be inserted into the image without affecting the main content and safety of the image and the location that the object should be inserted; (b) Inserting texts: Return appropriate texts that can be inserted into the image without affecting semantics of the image. (c) Style transfer: Choose a suitable style from {Watercolor style, Sketch style, Comic style} based on the image. Then we conduct manipulation operations by utilizing the ICEdit (Zhang2025InContextEE) model. Examples are presented in Figure 4.
@@ 3.4 Text-Image Dynamics
Section 3.2 introduces a set of text dynamic strategies. In this text-to-image strategy, we aim to inject text perturbations into images via cross-modal generation, thereby impacting MLLM safety. This strategy amplifies the influence of text dynamics through cross-modal interactions. Specifically, we first sample a from text dynamics, then we feed the sampled and original image into GPT-4o to generate an image caption and extract the safety-related keywords, which will then be used as a prompt to generate new images via stable diffusion (Esser2024ScalingRF). We then utilize the sampled text and the generated image as the new text-image pair.
@@
We mainly investigate two jailbreaking tricks. a) Figstep (Gong2023FigStepJL) proves that MLLMs can recognize and answer the typographic questions in images. The safety guardrails of MLLMs are ineffective against the typographic visual prompts. Even if the LLM part of MLLMs has been safety aligned in advance, the visual inputs could introduce new risks since the visual embedding space is not safety aligned to the LLM’s embedding space. Based on this, we directly replace the text prompts with their typographic version. Thus, in this trick, only the typographic prompt and the original image are fed into MLLMs. b) HADES (Li2024ImagesAA) empirically validates that MLLMs can be significantly affected by the unsafe words contained in images rather than texts, so we utilize GPT-4o to extract key information, which is strongly related to the safety content, and then inject it into the original image. This trick moves the unsafe content from the texts to the images without changing the original semantics.
@@ 4.1 Experimental Setup
In order to evaluate the effectiveness of SDEval, we select two comprehensive multimodal large language models safety benchmark MLLMGuard (gu2024mllmguard) and VLSBench (hu2024vlsbench) for dynamic evaluation.
MLLMguard (gu2024mllmguard) develops a collection of adversarial examples to test the ability of MLLMs to identify and counteract attacks orchestrated by red teams. We follow (gu2024mllmguard) to evaluate the proposed two metrics, Attack Success Degree (ASD), which quantifies the extent to which the responses are harmless, and Perfect Answer Rate (PAR), which represents the proportion of safe and responsible responses among all responses.
VLSBench (hu2024vlsbench) aims to address the existing issues in current multimodal safety benchmarks, called Visual Safety Information Leakage (VSIL). Following VLSBench, we calculate the Safety Rate (SR) by considering the total number of safe refusals and safe warnings.
@@
We evaluate four close source MLLMs: GPT-4o (gpt4o), o3 (openai2025o3), Claude-4-Sonnet (claude_4_opus), and Gemini2.5-Pro (Gemini2.5), and extensive open-sourced models: Qwen-VL family (qwen25vl): Qwen2.5-VL-7B, Qwen2.5-VL-72B, Qwen2-VL-7B, Qwen2-VL-72B; Yi-VL family (Young2024YiOF): Yi-VI-6B, Yi-VI-34B; InternVL family (internvl3): InternVL-3-9B, InternVL-3-14B, InternVL-Chat-V1-5; LLaVA family (liu2023llava): LLaVA-V1.5-7B, LLaVA-V1.5-13B. To ensure a standardized comparison, we set the generation temperature to 0 for all models. All experiments are conducted on 4 NVIDIA A800 GPUs.